How many molecules of water are there in 72g of H2O


1
Expert's answer
2020-09-14T08:23:37-0400

M (H2O) = 18 g/mol

n (H2O)=m/M=72/18=4 moles

Each mole of matter contains 6.022×1023 molecules.

Number of molecules in 72 g of water is:

6.022×1023 x 4 = 24.088 ×1023 molecules